Let one week's supply (S) and demand (D) functions for gasoline be given by p Upper D (q) equals 300 minus four-fifths q and p equals S(q) equals two-fifths q, where p is the price in dollars and q is the number of 42-gallon barrels. (A) On what interval is the quantity supplied below the equilibrium quantity?
Answer #1
Demand (D) = 300 - 4/5 q
Supply (S) = 2/5 q
Where, q = Equilibrium quantity
At Equilibrium point, Demand (D) = Supply (S)
Demand (D) = Supply (S)
300 - 4/5 q = 2/5 q
300 = 4/5 q + 2/5 q
300 = 6/5 q
q = (300*5) / 6
q = 250
Therefore, at quantity of 250 barrels.